OpenStudy (anonymous):

(secX+1)( secX-1)=(tan^2)X help me prove this using identies

OpenStudy (anonymous):

(a-b)(a+b)= a^2 - b^2 is an identity So: (secx - 1)(secx + 1) = sec^2(x) - 1 = 1/cos^2(x) - 1 = 1/cos^2(x) - cos^2(x)/cos^2(x) = [1 - cos^2(x)]/cos^2(x) = sin^2(x)/cos^2(x) = tan^2(x)